# Label Set Optimization via Activation Distribution Kurtosis for Zero-shot Classification with Generative Models

### v1

@@ Appendix B Data Leakage
As far as we know, Llama 3 and Llama 3.1 are not supervised fine-tuned with any public stance classification datasets. Flan-T5 is fine-tuned on Super-NaturalInstructions dataset (Wang et al., 2022), containing two English stance classification tasks (Kobbe et al., 2020) (i.e., task 209 and 513). The tasks are formed as a binary (“in favor” and “against”) and a three-way (“in favor”, “against” and “neutral”) classification, respectively. There is no overlapping between these two datasets and our nine experimented datasets.
@@ Appendix C Label Pool Creation
Following the pipeline we described in Section 2, we collect the seed label sets from the nine stance classification datasets (see Table 1). For the semeval dataset, the label set in the original paper ("favor, against, neither" in Table 1) is slightly different from the set used in their published dataset ("favor, against, none"), so we consider both of them.
@@ Appendix E Prompting with Label Explanation
More importantly, the benefits from label explanations in the prompt would not close the gap between the optimal and sub-optimal label sets, comparing the above improved performance with the maximum performances in Table 2.
@@ Appendix F Label Order Results
We supplement the maximum performance gain and drop on each dataset after re-ordering the label names for the top-k optimal and poor label sets with Llama 3.1 and Flan-T5-xl in Figure 4.
(a) Llama 3.1: top-k optimal label sets
(b) Llama 3.1: top-k poor label sets
(c) Flan-T5-xl: top-k optimal label sets
(d) Flan-T5-xl: top-k poor label sets
Figure 4: The maximum performance gain (positive value) and drop (negative value) on each dataset after re-ordering the label names for the top-k optimal and poor label sets with Llama 3.1 and Flan-T5-xl.
@@ Appendix G Label Elaboration Results
Llama 3.1 perspectrum 0.027 0.020 0.048 0.025 0.037 0.022
@@
Table 10: The average absolute performance change after elaborating for optimal or poor single-word label sets with Llama 3.1 and Flan-t5-xl ($E_{1}$, $E_{2}$, $E_{3}$ see Figure 1).
Appendix H Label Generalization Analysis
we report the overlap of top-5 optimal/poor ($L_{D_{i}\_M}\bigcap L_{D_{j}\_M}$) between each dataset on Llama 3, Llama 3.1 and Flan-T5-xl in Figure 5. The overlap of top-5 optimal/poor ($L_{D\_M_{i}}\bigcap L_{D\_M_{j}}$) between each model is presented in Figure 6.
(a) Llama 3 (Opt.)
(b) Llama 3 (Sub-Opt.)
(c) Llama 3.1 (Opt.)
(d) Llama 3.1 (Sub-Opt.)
(e) Flan-T5 (Opt.)
(f) Flan-T5 (Sub-Opt.)
Figure 5: Overlap among top-5 best (Opt.) or worst (Sub-Opt.) positive-negative stance label sets across datasets on Llama 3, Llama 3.1 and Flan-t5-xl respectively. Each entry $ij$ represents the number of overlapped best or worst label set between dataset on row $i$ and column $j$. The greater number indicate one label set is observed shared effectiveness across datasts.
(a) Best top5
(b) Worst top5
Figure 6: Overlap among top-5 best and worst positive-negative stance label sets on same dataset between different models. Each entry $ij$ represents the number of overlapped label sets for dataset $i$ (row $i$) between the models on column $j$. LLaMA3&LLaMA3.1: column $1$, LLaMA3&Flan-t5: column $2$, LLaMA3.1&Flan-t5: column $3$.
Appendix I Prompt Sensitivity Analysis of LOADS
@@
Due to the computational resource constraints, we manually craft two prompts and test LOADS with the four binary stance classification datasets on Llama 3. In the two prompts, we replace Given a [text1_name] and a [text2_name], detect the stance that the [text2_name] has towards the [text1_name] (see Section 3) with two different queries:
@@
As shown in Table 11, although different prompts with the same label sets may result in performance changes as expected (compare with Table 2), LOADS is robust to different prompts used for the label selection. The performance gap between LOADS-selected and original label sets tends to be similar across prompts.
Dataset Lowest Kurtosis Original Label
@@
Appendix J Perplexity Analysis
As discussed in Section 7, Gonen et al. (2023) empirically show that zero-shot ICL performance is statistically negative correlated with the perplexity of the prompt with input. However, they did not take into account the label options in the prompt when calculating the perplexity. Therefore, we further investigate whether the perplexity is also correlated with the variance zero-shot ICL performance caused by different label names.
Specifically, we use the prompt template in Section 3 and calculate the perplexity of prompts with inputs and different label sets. Following Gonen et al. (2023), for each label set, we average the perplexity over the dataset. And then we adopt spearman correlation test between the averaged perplexity scores and model performances. Due to the computational restriction, we experiment with all the binary datasets on Flan-T5-xl and Llama3-8b. Since Flan-T5 is an encoder-decoder model where perplexity has a loose definition, we treat the encoder input as an empty string when calculating perplexity.
@@
Appendix K Label Attention Key Similarity Analysis
In this section, we explore whether the closely related observation on few-shot ICL could be directly adopted to zero-shot ICL. Specifically, we focus on the study discussed in Section 7, where Wang et al. (2023) suggest that the when the LLM is prompted by demonstration with examples in a few-shot ICL setting, the model is likely to confuse the label categories if their key vectors in the attention modules are similar to each other.
@@
Appendix L Layer-wise Output Projections Analysis
We hypothesize that LLM may jump to the output prediction at last layers when a sub-optimal label set is used in the prompt. Therefore, we extract the hidden states from each decoder layer and project them on the model vocabulary, so that we obtain the ranked position of the final predicted label’s token in each layer Elhage et al. (2021); Geva et al. (2022)..
@@
Appendix M Human Translation Details

### v2

@@ Appendix B Data Leakage
As far as we know, Llama 3 and Llama 3.1 are not supervised fine-tuned with any public stance classification datasets. Flan-T5 is fine-tuned on Super-NaturalInstructions dataset Wang et al. (2022), containing two English stance classification tasks Kobbe et al. (2020) (i.e., task 209 and 513). The tasks are formed as a binary (“in favor” and “against”) and a three-way (“in favor”, “against” and “neutral”) classification, respectively. There is no overlapping between these two datasets and our nine experimented datasets.
@@ Appendix C Label Pool Creation
Stance Classification
Following the pipeline we described in Section 3 of the main paper, we collect the seed label sets from the nine stance classification datasets (see Table 1). For the semeval dataset, the label set in the original paper ("favor, against, neither" in Table 1 in main paper) is slightly different from the set used in their published dataset ("favor, against, none"), so we consider both of them.
@@
Topic Classification
Similarly, we follow the pipeline to collect and generate synonyms for each topic category. For TweetTopic dataset, since pop culture is a mixture of multiple sub-topics as discussed by Antypas et al.,Antypas et al. (2022), we also consider the synonyms of the sub-topics. We use every possible combination of synonyms among topic categories for TweetTopic dataset. For AG news, there are total 160 combinations. We randomly sample 50 of them due to limited computational resources. The number of label sets we experiment with for two datasets is listed in Table 6.
@@ Appendix E Prompting with Label Explanation
More importantly, the benefits from label explanations in the prompt would not close the gap between the optimal and sub-optimal label sets, comparing the above improved performance with the maximum performances in Table 2 in the main paper.
@@ Appendix F Label Order Results
The maximum performance gain and drop on each dataset after re-ordering the label names for the top-k optimal and poor label sets with Llama3, Llama 3.1 and Flan-T5-xl are in Figure 2.
(a) Llama3: top-k optimal label sets
(b) Llama3: top-k sub-optimal label sets
(c) Llama 3.1: top-k optimal label sets
(d) Llama 3.1: top-k sub-optimal label sets
(e) Flan-T5-xl: top-k optimal label sets
(f) Flan-T5-xl: top-k sub-optimal label sets
Figure 2: The maximum performance gain (positive value) and drop (negative value) on each dataset after re-ordering the label names for the top-k optimal and sub-optimal label sets with Llama3, Llama 3.1 and Flan-T5-xl.
@@ Appendix G Label Elaboration Results
Llama 3
perspectrum 0.016 0.018 0.010 0.015 0.017 0.009
ibmcs 0.027 0.041 0.024 0.014 0.029 0.044
snopes 0.055 0.040 0.054 0.026 0.029 0.018
emergent 0.051 0.053 0.047 0.038 0.022 0.040
rumoureval 0.095 0.036 0.084 0.020 0.058 0.102
arc 0.017 0.024 0.015 0.021 0.035 0.049
Llama 3.1
perspectrum 0.027 0.020 0.048 0.025 0.037 0.022
@@
Table 9: The average absolute performance change after elaborating for optimal or poor single-word label sets with Llama 3.1 and Flan-t5-xl ($E_{1}$, $E_{2}$, $E_{3}$ see Figure 1 in main paper).
Appendix H Prompt Sensitivity Analysis of LOADS
@@
Due to the computational resource constraints, we manually craft two prompts and test LOADS with the four binary stance classification datasets on Llama 3. In the two prompts, we replace Given a [text1_name] and a [text2_name], detect the stance that the [text2_name] has towards the [text1_name] (see Section 3.2 in main paper) with two different queries:
@@
As shown in Table 10, although different prompts with the same label sets may result in performance changes as expected (compare with Table 5 in main paper), LOADS is robust to different prompts used for the label selection. The performance gap between LOADS-selected and original label sets tends to be similar across prompt templates.
Dataset LOADS Original Label
@@
Appendix I Perplexity Analysis
As discussed in Section 2 in main paper, Gonen et al.,Gonen et al. (2023) empirically show that zero-shot ICL performance is statistically negative correlated with the perplexity of the prompt with input. However, they did not take into account the label options in the prompt when calculating the perplexity. Therefore, we further investigate whether the perplexity is also correlated with the variance zero-shot ICL performance caused by different label names.
Specifically, we use the prompt template in Section 3.2 in main paper, and calculate the perplexity of prompts with inputs and different label sets. Following Gonen et al.,Gonen et al. (2023), for each label set, we average the perplexity over the dataset. And then we adopt spearman correlation test between the averaged perplexity scores and model performances. Due to the computational restriction, we experiment with all the binary datasets on Flan-T5-xl and Llama3-8b. Since Flan-T5 is an encoder-decoder model where perplexity has a loose definition, we treat the encoder input as an empty string when calculating perplexity.
@@ Appendix I Perplexity Analysis
Appendix J Label Attention Key Similarity Analysis
In this section, we explore whether the closely related observation on few-shot ICL could be directly adopted to zero-shot ICL. Specifically, we focus on the study discussed in Section 2 in main paper, where Wang et al.,Wang et al. (2023) suggest that the when the LLM is prompted by demonstration with examples in a few-shot ICL setting, the model is likely to confuse the label categories if their key vectors in the attention modules are similar to each other.
@@
Appendix K Layer-wise Output Projections Analysis
We hypothesize that LLM may jump to the output prediction at last layers when a sub-optimal label set is used in the prompt. Therefore, we extract the hidden states from each decoder layer and project them on the model vocabulary, so that we obtain the ranked position of the final predicted label’s token in each layer Elhage et al. (2021); Geva et al. (2022).
@@
Appendix L Human Translation Details